OpenStudy (anonymous):

What is the equation of the line passing through the points (2, –1) and (5, –10) in slope-intercept form? y=-3x-5 y=-3x+5 y=3x-5 y=3x+5

OpenStudy (anonymous):

y = mx+b find m first: \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] then plug either (x,y) points given in question along with m to find what b is

OpenStudy (anonymous):

still confused:(

OpenStudy (anonymous):

do u just want the answer or do you want to know how?

OpenStudy (anonymous):

the answer im taking a test and i reallly need a good grade :(

OpenStudy (anonymous):

y = -3x + 5

OpenStudy (anonymous):

can you help me with more plzzz ill give u alot of medals :(

OpenStudy (anonymous):

most u can give is one lol. but i guess i can help

OpenStudy (anonymous):

aww thank you :)

OpenStudy (anonymous):

What is the equation of the line passing through the points (–25, 50) and (25, 50) in slope-intercept form?

OpenStudy (anonymous):

y = 50

OpenStudy (anonymous):

y=50x y=-50 y=50x y=50

OpenStudy (anonymous):

y=-50*

OpenStudy (anonymous):

y = 50

OpenStudy (anonymous):

:) i got more

OpenStudy (anonymous):

lets get it over with lol

OpenStudy (anonymous):

Which is the equation of a line that has a slope of 1 and passes through point (5, 3)?

OpenStudy (anonymous):

y = x - 2

OpenStudy (anonymous):

The point-slope form of a line that has a slope of –2 and passes through point (5, –2) is shown below.

OpenStudy (anonymous):

y = -2x + 8

OpenStudy (anonymous):

y+2=-2(x-5)

OpenStudy (anonymous):

that

OpenStudy (anonymous):

they want the slope form

OpenStudy (anonymous):

slope intercept form

OpenStudy (anonymous):

gave it: y = -2x + 8

OpenStudy (anonymous):

ok ok lol sorry :(

OpenStudy (anonymous):

lol its np

OpenStudy (anonymous):

Which is a characteristic of the line that passes through the points (6, 10) and (12, 7)?

OpenStudy (anonymous):

the slope is 1/2 the slope is -2 the y intercept is 7 the y intercept is 13

OpenStudy (anonymous):

y intercept is 13

OpenStudy (anonymous):

Leon correctly found the slope and y-intercept of the line that passes through the points (9, –8) and (3, 4) as follows.

OpenStudy (anonymous):

What is the equation of the line in slope-intercept form?

OpenStudy (anonymous):

y = -2x + 10

OpenStudy (anonymous):

lol they show you exactly how to do it. you should keep that image to learn from

OpenStudy (anonymous):

Mike correctly found the slope and y-intercept of the line passing through the points (–5, –2) and (3, 14) as follows. What is the equation of the line in slope-intercept form?

OpenStudy (anonymous):

y = 2x + 8

OpenStudy (anonymous):

thank you :) sooo much

OpenStudy (anonymous):

:)